Exercise 16-47 (Algo) Overhead Variances (LO 16-5, 6)

The (partial) cost sheet for the single product manufactured at Vienna Company follows.

 






Direct labor(17 hours @ $25)$425
Variable overhead(17 hours @ $2)
34
Fixed overhead(17 hours @ $4)
68

 
The master budget level of production is 82,000 direct-labor hours, which is also the production volume used to compute the fixed overhead application rate. Other information available for operations over the past accounting period include the following.
 





Actual variable overhead incurred$132,000
Actual fixed overhead incurred
343,200
Direct labor efficiency variance
162,000U
Variable overhead price variance
30,000F

 

Required:
a. What was the variable overhead efficiency variance?
b. What was the fixed overhead price variance?
c. What was the fixed overhead production volume variance?
Answer #1

A) step 1) 30,000= (2* actual hours)- 132,000

actual hours= 132,000+ 30,000/ 2= 147,000

Step 2) -162,000= (standard hours* 25)- (147,000*25)

  3,675,000- 162,000/ 25= 12,960 U answer Variable overhead efficiency variance


B) (82,000* 4)- 343,200= 15,200 U Fixed overhead price variance


C) 

step 1) 2*82,000= 164,000

step 2) 25* 82,000= 2,050,000

step 3) 162,000/ 25= 6,480

2,050,000+162,000= 2,212,000

 2,212,000/ 25= 88,480

88,480* 2= 176,960

176,960- 164,000= 12,960 Answer for A


4*82,000= 328,000

323,200-328,000= 15,200 answer for B


132,000+ 30,000= 162,000

162,000/2= 81,000

81,000- 6,480= 74,520

4* 74,520= 298,080

328,000(part B)- 298,080= 29,920 fixed overhead production volume variance answer